# Separating subsets from their images

### v1

@@ Abstract.
Let $G$ be a transitive permutation group. In this paper, we introduce and study the parameter $\mathbf{m}(G)$, which estimates the size of the largest set of points $A$ such that there exists a permutation $g$ with the property that $A\cap A^{g}$ is empty. In particular, we focus on deriving general bounds for arbitrary transitive groups, and on the asymptotic behaviour of certain families of primitive groups.
@@
The research presented in this paper was initiated during a research visit by MB and KR to IMFM, funded by Slovenian Research and Innovation Agency, research programme number P1-0294. They gratefully acknowledge the support and hospitality provided. MB is supported by the Slovenian Research Agency program P1-0222 and grant J1-50001, and he is a member of the Italian GNSAGA INdAM research group. ML and $P$ are supported by the Slovenian Research and Innovation Agency, programme number P1-0294.
@@ 1. Introduction
Let us begin by presenting a graph-theoretical problem [^1] We thank Urban Jezernik for pointing this problem to us. which spurred our interest in the topic of the paper and led us to define and investigate the notions presented here.
Suppose we are given a simple finite graph $\Gamma$ on $n$ vertices and we want to know whether its complement contains an isomorphic copy of $\Gamma$ as a subgraph, or equivalently, whether two copies of $\Gamma$ can be packed into the complete graph $\mathbf{K}_{n}$. Clearly, if $\Gamma$ contains more than $n\choose 2$} edges, then such a packing is impossible. On the other hand, if $\Gamma$ is very sparse (for example, if it has maximum valence $1$ and at least $3$ vertices), then the packing will very likely exist. It is thus natural to ask what is the largest integer $m$ such that every graph on $n$ vertices and at most $m$ edges can be packed twice into $\mathbf{K}_{n}$. Since such a packing is clearly impossible for the complete bipartite graph $\mathbf{K}_{1,n-1}$, we see that $m\leq n-2$. Surprisingly, as was ingeniously proved in [12, Theorem 1], this bound is sharp, in the sense that every graph on $n$ vertices and at most n2n-2 edges lies in its complement as a subgraph.
What we just discussed for graphs can actually be translated into the language of permutation groups. Suppose $\Gamma$ is a graph with vertex-set $n]:=\{1,2,\ldots,n\}$ and let $E\subseteq{[n]\choose 2}$ be its edge-set. If there exists an isomorphism from $\Gamma$ to a graph contained in the complement of $\Gamma$, then this isomorphism is a permutation of $[n]$ which, when viewed as a permutation on the unordered pairs $n]\choose 2$}, maps $E$ to a set disjoint from $E$. In short, two copies of $\Gamma$ can be packed into $\mathbf{K}_{n}$ if, and only if, the group $\mathrm{Sym}(n)$ in its action on $n]\choose 2$} contains an element $g$ such that $E\cap E^{g}=\emptyset$; we shall say that such a set $E$ is self-separable—see Definition 1. The graph-theoretical theorem mentioned in the previous paragraph can now be translated into the assertion that the size of a smallest non-self-separable set for the action of $\mathrm{Sym}(n)$ on $n]\choose 2$} is n1n-1.
As we will try to show in this paper, putting the original packing problem in this general group-theoretical context reveals connections to several classical problems and theorems in combinatorics, finite geometry and group theory, as well as opens a plethora of new and interesting questions, some of which will be addressed in this paper.
@@
Observe that the parameter $\mathbf{m}(G)$ is also equal to the largest integer such that all sets of size less than that integer are self-separable, that is:
@@
The above question can of course be approached from several angles, ranging from algorithmic considerations to exact theoretical results.
The computational task of determining $\mathbf{m}(G)$ for a given permutation group $G$ appears to be highly challenging. The only algorithm we devised during the preparation of this paper involves an exhaustive search (see [5] and Section 7): Systematically testing all subsets $A$ of the domain whose sizes exceed the theoretical lower bound we will establish in C, and stopping when we find the minimal example $A$ that is not self-separable for $G$. When $G$ is known to be $k$-homogenous, we may limit the search space to subset containing a fixed $k$-subset. No further improvements of this naïve approach are known to us. Therefore we ask the following:
@@
In this paper, however, we shall be especially interested in asymptotic (lower and upper) bounds on $\mathbf{m}(G)$ in terms of the degree of $G$. As we shall see, for every group $G$ of order $n$, the parameter $\mathbf{m}(G)$ always satisfies $\sqrt{n}\leq\mathbf{m}(G)\leq\lceil(n+1)/2\rceil$ (see C and Example 5). Moreover, infinite families of transitive groups exist for which $\mathbf{m}(G)$ is asymptotically square root of the degree (for example, the family of symmetric groups acting on unordered pairs) as well as infinite families where $\mathbf{m}(G)$ is asymptotically linear in the degree of the group (for example, the symmetric groups in their natural action). What we find interesting, though, is that there seems to be a dichotomy between these square root and linear asymptotical behaviours.
Let us present the remaing players of this paper. For an infinite family $\mathcal{C}$ of finite permutation groups and for an integer $n$, let $\mathcal{C}_{n}$ be the subfamily of permutation groups in $\mathcal{C}$ that have degree $n$, and let
@@
To understand the behavior of these functions on large class of transitive permutation groups, we shall tackle the following problem.
@@
To give you the taste of the statements we will be proving, let us list some of them. In the case of the class $\mathcal{C}_{{\rm tr}}$ of all transitive permutation groups, we prove that $\liminf\mathbf{f}_{\mathcal{C}_{{\rm tr}}}(n)$ grows asymptotically as $\sqrt{n}$ (see Theorem D), while $\limsup\mathbf{F}_{\mathcal{C}_{{\rm tr}}}(n)$ is asymptotically $\frac{n}{2}$ (see Theorem C), or to be precise:
@@
Further, as is stated in Theorem E, for the class $\mathcal{C}_{{\rm reg}}$ of regular permutation groups, the asymptotic bound for $\mathbf{F}_{\mathcal{C}_{{\rm reg}}}$ can be significantly improved to
@@
A particular attention is also given to the class $\mathcal{C}_{{\rm pr}}$ of primitive permutation groups, where we start to address the following problem.
@@ 2.1. General bounds
Let $G$ be a transitive permutation group of degree $n$, and let $\alpha$ be a point. Then
@@
with equality if, and only if, $G$ is a regular group of collineations of a projective plane. In particular, if $\mathcal{C_{\mathrm{tr}}}$ is the class of all transitive permutation groups, then
@@
The proof of C (given in section Section 4) relies on the Neumann Separation Lemma (see [8, 37]), and the characterisation of the equality depends on considerations from incidence geometry.
@@
The following example shows that the bound is sharp.
Example 5.
Let $G$ be either the symmetric group $\mathrm{Sym}(n)$ or the alternating group $\mathrm{Alt}(n)$ endowed with their natural actions on $[n]$. Since $G$ is $\lfloor n/2\rfloor$-transitive, every subset containing half of the points or less can be mapped to its complement by a permutation in $G$. Hence, every subset $A$ containing $\lfloor n/2\rfloor$ points is self-separable for $G$. Therefore,
$\mathbf{m}(\mathrm{Sym}(n))=\mathbf{m}(\mathrm{Alt}(n))=\left\lfloor\frac{n}{2}\right\rfloor+1=\left\lceil\frac{n+1}{2}\right\rceil.$
Therefore, we obtain the following asymptotic behavior.
@@
Classifying the groups that attain the upper bound in Inequality (B) is a nontrivial problem, posing challenges both computationally and theoretically. We provide this classification in Section 7. Broadly speaking, unless the degree of $G$ is less than $24$, $G$ is either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ in their natural actions on $n$ points, or it is isomorphic to a subgroup of $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(m)$ or $\mathrm{Sym}(m)\mathop{\mathrm{wr}}C_{2}$ endowed with the imprimitive action on $2m$ points (where the permutation group that $G$ induces on the system of blocks or on a single block, respectively, is either alternating or symmetric). For a precise statement, including the complete list of small-degree groups meeting the bound, we refer to Lemmas 21 and 22.
2.2. Permutation groups with bounded stabiliser
In the previous section, we determined the asymptotic behaviour of $\mathbf{f}_{\mathcal{C}_{\mathrm{tr}}}$ and $\mathbf{F}_{\mathcal{C}_{\mathrm{tr}}}$ for the class of all transitive permutation groups. Note however, that the gap between the square root asymptotics of $\liminf\mathbf{f}_{\mathcal{C}_{\mathrm{tr}}}(n)$ and the linear growth of $\mathbf{F}_{\mathcal{C}_{\mathrm{tr}}}(n)$ leaves many questions about the parameter $\mathbf{m}(G)$ for a generic transitive permutation group $G$ open. In what follows, we will try to identify subclasses $\mathcal{C}\subseteq\mathcal{C}_{\mathrm{tr}}$ for which we are able to exert a more effective control. That is, families for which the gap between $f_{\mathcal{C}_{\mathrm{tr}}}$ and $F_{\mathcal{C}_{\mathrm{tr}}}$ is smaller. In fact, we will observe that for many important families of permutation group $\mathcal{C}\subseteq\mathcal{C}_{\mathrm{tr}}$, $\mathbf{F}_{\mathcal{C}}(n)$ is asymptotically proportional to $\sqrt{n}$ (as well as $\mathbf{f}_{\mathcal{C}}(n)$, by consequence). The class of regular permutation groups provides a first example supporting this viewpoint.
Theorem E.
@@ 2.2. Permutation groups with bounded stabiliser
This result is essentially proven in Remark 18, where it is shown that, for a subset $A$ not being self-separable for a regular group $G$ is equivalent to $A$ being a difference basis for $G$:
@@
Let us remark that the asymptotics for the minimal size of a difference basis for a generic group $G$ is still an open problem in additive combinatorics.
If every nontrivial element of $G$ can be written uniquely as $ab^{-1}$ for some $a,b\in A$, then the difference basis $A$ is called a difference set, which are a widely used tool in finite geometries. If $q$ is the order of a classical projective plane, the celebrated paper [45] by Singer proves the existence of a difference set of size $q+1$ for the cyclic group of order $q^{2}+q+1$. Recall that these permutation groups are precisely those that meet the lower bound in C.
@@ 2.3. Reduction to primitive permutation groups
Recall that every imprimitive permutation group $G$ can be reduced to an iterated wreath product of primitive permutation groups as follows. Let $\Sigma$ be a system of imprimitivity for $G$ and let $B$ be a block in $\Sigma$. As is standard, we denote by $G_{B}^{B}$ the permutation group that the setwise stabiliser $G_{B}$ induces on $B$, and by $G^{\Sigma}$ the permutation group that $G$ induces on $\Sigma$. Then $G$ embeds into the imprimitive wreath product $G_{B}^{B}\mathop{\mathrm{wr}}G^{\Sigma}$. Moreover, if $\Sigma$ is a coarsest system of imprimitivity, then $G^{\Sigma}$ is primitive, meanwhile if $\Sigma$ is a finest system of imprimitivity, then $G_{B}^{B}$ is primitive.
@@
Theorem H.
Let $G$ be a group, and suppose that $G$ embeds into the imprimitive wreath product $K\mathop{\mathrm{wr}}H$. Then
@@
The upper bound is asymptotically sharp. Let $K$ and $H$ be isomorphic to the cyclic group of degree $q^{2}+q+1$ with regular action, and let $G=K\mathop{\mathrm{wr}}H$. Note that the degree of $G$ is $(q^{2}+q+1)^{2}$. For large $q$, using C both to find the lower bound for $\mathbf{m}(G)$ and the upper bounds for $\mathbf{m}(K)$ and $\mathbf{m}(H)$,
@@
This proves the asymptotic sharpness of the upper bound of H. The same reasoning, applied with $K$ and $H$ taken as regular groups and using E, shows that $\mathbf{m}(G)\leq\frac{8}{3}\sqrt{n}$ – matching the optimal lower bound up to a multiplicative constant.
On the other hand, we are unable to prove the optimality of the lower bound. Even showing the existence of a family of permutation groups for which the lower bound is asymptotically tight up to a constant requires further work. The following ingredient, of independent interest, will be essential. (Its proof is given in Section 6.)
Theorem I.
@@
Let $a$ and $b$ be two even positive integers, and let $K=\mathrm{Sym}(a)$ and $H=\mathrm{Sym}(b)$. By choosing $G=K\mathop{\mathrm{wr}}H$, I implies
@@
On the other hand, H yields
@@
If we allow $a$ and $b$ to grow as two arbitrary unbounded functions, we find that, asymptotically, $\mathbf{m}(G)$ is twice the lower bound of H – proving the asymptotic sharpness up to the multiplicative constant $2$. This leads to the following question.

### v2

@@ Abstract.
Let $G$ be a transitive permutation group acting on $\Omega$. In this paper, we introduce and study the parameter $\mathbf{m}(G)$, which denotes the size of the smallest set of points $A$ such that, for every permutation $g\in G$, $A\cap A^{g}$ is nonempty. In particular, we focus on deriving general bounds for arbitrary transitive groups, and on the asymptotic behaviour of certain families of primitive groups. We also provide a classification of transitive groups with $\mathbf{m}(G)$ largest possible, namely with $\mathbf{m}(G)=\lceil(|\Omega|+1)/2\rceil$.
@@
The research presented in this paper was initiated during a research visit by MB and KR to IMFM, funded by Slovenian Research and Innovation Agency, research programme number P1-0294. They gratefully acknowledge the support and hospitality provided. MB is supported by the Slovenian Research Agency programme P1-0222 and grant J1-50001, and he is a member of the Italian GNSAGA INdAM research group. ML and $P$ are supported by the Slovenian Research and Innovation Agency, programme number P1-0294.
@@ 1. Introduction
Let us begin by presenting a graph-theoretical problem which spurred our interest in the topic of the paper and led us to define and investigate the notions presented here.
Suppose we are given a simple finite graph $\Gamma$ on $n$ vertices and we want to know whether its complement contains an isomorphic copy of $\Gamma$ as a subgraph, or equivalently, whether two copies of $\Gamma$ can be packed into the complete graph $\mathbf{K}_{n}$. Clearly, if $\Gamma$ contains more than $2^{-1}{n\choose 2}$ edges, then such a packing is impossible. On the other hand, if $\Gamma$ is very sparse (for example, if it has maximum valence $1$ and at least $3$ vertices), then the packing will very likely exist. It is thus natural to ask what is the largest integer $m$ such that every graph on $n$ vertices and at most $m$ edges can be packed twice into $\mathbf{K}_{n}$. Since such a packing is clearly impossible for the complete bipartite graph $\mathbf{K}_{1,n-1}$, we see that $m\leq n-2$. Surprisingly, as was elegantly proved in [12, Theorem 1], this bound is sharp, in the sense that every graph on $n$ vertices and at most n2n-2 edges lies in its complement as a subgraph.
The above problem for graphs can actually be translated into the language of permutation groups. Suppose $\Gamma$ is a graph with vertex set $n]:=\{1,2,\ldots,n\}$ and edge set $E\subseteq{[n]\choose 2}$. An isomorphism from $\Gamma$ to a graph contained in the complement of $\Gamma$, is a permutation of $[n]$ which, when viewed as a permutation on the unordered pairs $n]\choose 2$}, maps $E$ to a set disjoint from $E$. In short, two copies of $\Gamma$ can be packed into $\mathbf{K}_{n}$ if and only if the group $\mathrm{Sym}(n)$ in its action on $n]\choose 2$} contains an element $g$ such that $E\cap E^{g}=\emptyset$. We say that such a set $E$ is self-separable, and if there is no such element $g$ then we say it is non-self-separable (see Definition 1). The graph-theoretical theorem mentioned above is equivalent to the assertion that the size of the smallest non-self-separable set for the action of $\mathrm{Sym}(n)$ on $n]\choose 2$} is n1n-1.
In this paper we show that viewing the original packing problem in this general group-theoretical context reveals connections to several classical problems and theorems in combinatorics, finite geometry and group theory, as well as opens a plethora of new and interesting questions. We will address some of these and suggest further directions to explore.
@@
Observe that the parameter $\mathbf{m}(G)$ is also equal to the largest integer such that all sets of size less than that integer are self-separable, that is,
@@
The above question can be approached from several angles, ranging from algorithmic considerations to exact theoretical results. We mainly focus on theoretical ways, but briefly consider the computational approach too.
The computational task of determining $\mathbf{m}(G)$ for a given permutation group $G$ appears to be highly challenging. The only algorithm we devised during the preparation of this paper involves an exhaustive search (see [5] and Section 7) using a theoretical lower bound we establish in C. Our algorithm systematically tests all subsets $A$ of the domain whose sizes exceed the theoretical lower bound in C, and stops when we find the minimal example $A$ that is non-self-separable for $G$. When $G$ is known to be $k$-homogeneous, we may limit the search space to subset containing a fixed $k$-subset. No further improvements of this naïve approach are known to us. Therefore we ask the following:
@@
In this paper, however, we will be especially interested in asymptotic (lower and upper) bounds on $\mathbf{m}(G)$ in terms of the degree of $G$.  C tells us that, for every group $G$ of order $n$, the parameter $\mathbf{m}(G)$ always satisfies $\sqrt{n}\leq\mathbf{m}(G)\leq\lceil(n+1)/2\rceil$. Moreover, infinite families of transitive groups exist for which $\mathbf{m}(G)$ is asymptotically square root of the degree (for example, the family of symmetric groups acting on unordered pairs, which is our motivating example) as well as infinite families where $\mathbf{m}(G)$ is asymptotically linear in the degree of the group (for example, the symmetric groups in their natural action).
To understand asymptotic behaviours better, we need the following definitions. For an infinite family $\mathcal{C}$ of finite permutation groups and for an integer $n$, let $\mathcal{C}_{n}$ be the subfamily of permutation groups in $\mathcal{C}$ that have degree $n$, and let
@@
To understand the behaviour of these functions on large class of transitive permutation groups, we will consider the following problem.
@@
To give a taste of the statements that we will be proving, let us list some of them. In the case of the class $\mathcal{C}_{{\rm tr}}$ of all transitive permutation groups, we prove that $\liminf\mathbf{f}_{\mathcal{C}_{{\rm tr}}}(n)$ grows asymptotically as $\sqrt{n}$ (see Theorem E), while $\limsup\mathbf{F}_{\mathcal{C}_{{\rm tr}}}(n)$ is asymptotically $\frac{n}{2}$ (see Theorem C), or to be precise:
@@
Further, as is stated in Theorem F, for the class $\mathcal{C}_{{\rm reg}}$ of regular permutation groups, the asymptotic bound for $\mathbf{F}_{\mathcal{C}_{{\rm reg}}}$ can be significantly improved to
@@
Particular attention is also given to the class $\mathcal{C}_{{\rm pr}}$ of primitive permutation groups, specifically to the following problem.
@@
1.1. Acknowledgements
We thank Urban Jezernik for drawing our attention to the graph-theoretic problem that motivated this work. We are also grateful to Pablo Spiga for several helpful conversations, particularly regarding Neumann’s Separation Lemma, and for identifying an error in an earlier version of the proof of Lemma 20.
@@ 2.1. General bounds
Let $G$ be a transitive permutation group on $\Omega$ with $|\Omega|=n$, and let $\alpha\in\Omega$ be a point. Then
@@
with equality if and only if $G$ is a regular group of collineations of a projective plane. In particular, if $\mathcal{C_{\mathrm{tr}}}$ is the class of all transitive permutation groups, then
@@
The proof of C (given in Section 4) relies on the Neumann Separation Lemma (see [8, 34]), and the characterisation of the equality depends on considerations from incidence geometry.
@@
Classifying the groups that attain the upper bound in Inequality (B) is a nontrivial problem and is interesting in its own right. For example, a very natural subproblem that needs to be addressed is: “Determine all primitive permutation groups $G$ of even degree such that every subset of the domain of size half of the degree can be mapped to its complement by an element of GG.” Clearly, the alternating and symmetric groups have this property, but it is somewhat surprising that so do some other permutation groups.
@@
Let $G$ be a transitive permutation group of degree $n$. If $G$ attains the upper bound in Inequality (B), then $G$ satisfies one of the following conditions.
  1. $(a)$
    $G$ is primitive, and either $n\leq 24$ or $\mathrm{Alt}(\Omega)\leq G\leq\mathrm{Sym}(\Omega)$. (The precise list of such groups can be found in Lemma 20.)
  2. $(b)$
    $G$ is imprimitive with a system of imprimitivity consisting of $n/2$ blocks of size $2$, and either $n\leq 24$ or the action on the block system is isomorphic to either $\mathrm{Alt}(n/2)$ or $\mathrm{Sym}(n/2)$. (The precise list of such groups can be found in Lemma 31.)
  3. $(c)$
    $G$ is imprimitive with a system of imprimitivity consisting of two blocks of size $n/2$, and either $n\leq 24$ or the action induced on a block is isomorphic to either $\mathrm{Alt}(n/2)$ or $\mathrm{Sym}(n/2)$. (The precise list of such groups can be found in Lemma 37.)
We stress that not all groups with the conditions above reach the putative upper bound. A complete characterisation is provided in Section 7 (once again, see Lemmas 20, 31 and 37).
Since D yields infinite families of permutation groups attaining the bound in Inequality (B), we have the following consequence:
Theorem E.
@@ 3. $(c)$
2.2. Permutation groups with bounded stabilisers
In the previous section, we determined the asymptotic behaviour of $\mathbf{f}_{\mathcal{C}_{\mathrm{tr}}}$ and $\mathbf{F}_{\mathcal{C}_{\mathrm{tr}}}$ for the class of all transitive permutation groups. Note, however, that the gap between the square root asymptotics of $\liminf\mathbf{f}_{\mathcal{C}_{\mathrm{tr}}}(n)$ and the linear growth of $\mathbf{F}_{\mathcal{C}_{\mathrm{tr}}}(n)$ leaves many questions about the parameter $\mathbf{m}(G)$ for a generic transitive permutation group $G$ open. In what follows, we will try to identify subclasses $\mathcal{C}\subseteq\mathcal{C}_{\mathrm{tr}}$ for which we are able to exert a more effective control. That is, families for which the gap between $f_{\mathcal{C}_{\mathrm{tr}}}$ and $F_{\mathcal{C}_{\mathrm{tr}}}$ is smaller. In fact, we will observe that for many important families of permutation group $\mathcal{C}\subseteq\mathcal{C}_{\mathrm{tr}}$, $\mathbf{F}_{\mathcal{C}}(n)$ is asymptotically proportional to $\sqrt{n}$ (as well as $\mathbf{f}_{\mathcal{C}}(n)$, by consequence). The class of regular permutation groups provides such a family.
Theorem F.
@@ 2.2. Permutation groups with bounded stabilisers
We will prove this result in Section 5. This result follows from Remark 17, which asserts that for a subset $A$ not being self-separable for a regular group $G$ is equivalent to $A$ being a difference basis for $G$:
@@
Let us remark that asymptotic results for the minimal size of a difference basis for a generic group $G$ is still an open problem in additive combinatorics.
If every nontrivial element of $G$ can be written uniquely as $ab^{-1}$ for some $a,b\in A$, then the difference basis $A$ is called a difference set, which is a widely used tool in finite geometries. For instance, if $q$ is the order of a classical projective plane, a paper [43] by Singer proves the existence of a difference set of size $q+1$ for the cyclic group of order $q^{2}+q+1$. Observe that these permutation groups are precisely those that meet the lower bound in C.
@@ 2.3. Reduction to primitive permutation groups
Recall that every imprimitive permutation group $G$ can be reduced to an iterated wreath product of primitive permutation groups as follows. Let $\Sigma$ be a system of imprimitivity for $G$ and let $B$ be a block in $\Sigma$. As usual, we denote the permutation group that the setwise stabiliser $G_{B}$ induces on $B$ by $G_{B}^{B}$, and the permutation group that $G$ induces on $\Sigma$ by $G^{\Sigma}$. Then $G$ embeds into the imprimitive wreath product $G_{B}^{B}\mathop{\mathrm{wr}}G^{\Sigma}$. Moreover, if $\Sigma$ is a coarsest system of imprimitivity, then $G^{\Sigma}$ is primitive, meanwhile if $\Sigma$ is a finest system of imprimitivity, then $G_{B}^{B}$ is primitive.
@@
Theorem I.
Let $G$ be a group, and suppose that $G$ embeds into the imprimitive wreath product $H\mathop{\mathrm{wr}}K$. Then
@@
The following example shows that the upper bound is asymptotically sharp. Let $K$ and $H$ be isomorphic to the cyclic group of degree $q^{2}+q+1$ with regular action, and let $G=K\mathop{\mathrm{wr}}H$. Note that the degree of $G$ is $(q^{2}+q+1)^{2}$. For large $q$, using C to find the lower bound for $\mathbf{m}(G)$ and the upper bounds for $\mathbf{m}(K)$ and $\mathbf{m}(H)$,
@@
This proves the asymptotic sharpness of the upper bound of I.
Furthermore, if $K$ and $H$ are regular groups, using the above argument and F, we can show that $\mathbf{m}(G)\leq\frac{8}{3}\sqrt{n}$ – matching the optimal lower bound up to a multiplicative constant.
On the other hand, we are unable to prove the sharpness of the lower bound. We show, however, that there exists a family of permutation groups for which the lower bound is asymptotically tight up to a constant. The following result, of independent interest, will be essential in finding such a family.
Theorem J.
@@
The proof of this result is given in Section 6. Let $a$ and $b$ be two even positive integers, and let $K=\mathrm{Sym}(a)$ and $H=\mathrm{Sym}(b)$. By choosing $G=K\mathop{\mathrm{wr}}H$, J implies that
@@
On the other hand, I yields
@@
If we allow $a$ and $b$ to grow as two arbitrary unbounded functions, we find that, asymptotically, $\mathbf{m}(G)$ is twice the lower bound of I – proving the asymptotic sharpness up to the multiplicative constant $2$. This leads to the following questions.
Question 6.
Is the lower bound stated in I asymptotically sharp? If it is not, is $2(\mathbf{m}(K)+\mathbf{m}(H))$ an asymptotically sharp lower bound? Furthermore, by excluding an appropriate infinite family of transitive groups, is it possible to obtain a larger lower bound?
2.4. Primitive permutation groups
Recall that, for $\mathrm{Alt}(n)$ and $\mathrm{Sym}(n)$ in their action on $n$ points, $\mathbf{m}(\mathrm{Alt}(n))=\mathbf{m}(\mathrm{Sym}(n))=\lceil(n+1)/2\rceil$. At the moment, they constitute the single infinite family of primitive permutation groups where we can prove that a linear asymptotic behaviour is exhibited. We suspect that, possibly apart from few well-understood infinite families, the size of the smallest non-self-separable subset is asymptotically proportional to the square root of the degree of the group. This intuition is captured by the following question.